# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v2

@@
Rezvaneh Rezapour6, Mary Ellen Sloane7, and Alexandre Bovet 1,2
@@
University of Zurich, 8057 Zurich, Swizterland
@@
Yasaman.asgari@math.uzh.ch
The Arab Spring is a major socio-political movement that reshaped democratic aspirations in the Middle East and North Africa, attracting global attention through news, social media, and academic discourse. However, its consequences on the academic landscape in the region are still unclear. Here, we conduct the first study of scholarly attention toward 10 target countries affected by the Arab Spring by analyzing more than 25 million articles published from 2002 to 2019. Using a difference-in-difference statistical framework, we find that most target countries have experienced a significant increase in scholarly attention post-Arab Spring compared to the rest of the world, with Egypt attracting the most attention. We investigate how funding and migration networks relate to scholarly attention and reveal that Saudi Arabia has emerged as a key player among Western nations by attracting researchers and funding projects that shape research on the region.
Introduction
Social upheavals like political revolutions and armed conflicts profoundly disrupt scientific ecosystems, precipitating research productivity decline, diminished international collaboration, and significant brain drain  (?, ?, ?, ?, ?, ?). The “Arab Spring” exemplifies this phenomenon—a transformative regional movement that started in Tunisia in 2010, where Mohamed Bouazizi’s self-immolation catalyzed widespread protests against autocratic regimes and systemic economic inequities  (?). This movement has had significant effects, from individual hardships, such as personal losses, displacement, and economic difficulties, to broader societal impacts, including social fragmentation, political and economic crises, and altered demographics as governments reformed or collapsed  (?). These widespread changes also significantly influenced global geopolitics  (?), migration patterns  (?), and research activities  (?).
In particular, the Arab Spring reshaped the political landscape across the Middle Eastern and North African (MENA) region. Its impact varied across countries, broadly grouping affected nations into five categories  (?): (1) countries where governments were overthrown, such as Egypt and Tunisia  (?, ?), (2) countries that descended into civil wars, such as Libya, Syria, and Yemen  (?, ?, ?), (3) countries that experienced significant protests leading to governmental changes, such as Jordan, Kuwait, Morocco, Bahrain, and Oman  (?, ?), (4) countries with major protests, such as Iraq, Lebanon, and Algeria, and (5) countries with minor protests, including Saudi Arabia, Somalia, Sudan, Palestine, Djibouti, and Mauritania. These differing trajectories influenced political and economic structures and academic environments  (?, ?), shaping research output  (?, ?, ?, ?, ?, ?), collaboration networks  (?, ?, ?), and scientific mobility  (?) in distinct ways.
One significant consequence of the Arab Spring was the surge in scholarly attention to the region. This refers to the extent to which research publications, conferences, and academic collaborations focus on specific countries or regions  (?, ?, ?). The increase was particularly evident in political science, international relations, and sociology, reflecting a heightened academic discourse on the MENA region’s transformations  (?, ?).
Beyond the Arab Spring’s impact on scholarly attention, broader structural disparities exist in global knowledge production. Historically, scientific attention and academic output have been concentrated in wealthier nations, while many regions—including MENA—have struggled with underrepresentation in global research networks, especially in health-related  (?, ?, ?, ?, ?, ?, ?, ?) topics.
In this context, distinguishing between domestic and foreign scholarly attention is essential for understanding the broader dynamics of academic engagement with a particular region  (?). Domestic scholarly attention encompasses research conducted by scholars on their own nation, often shaped by local priorities, resources, and institutional frameworks. On the other hand, foreign scholarly attention captures the perspectives and contributions of external researchers who engage with the region, often in response to significant events, policy shifts, or emerging research opportunities. The balance between these two forms of attention has significant implications for scientific productivity, local knowledge generation, and the region’s integration into global academic networks.
However, scholarly attention captures only one dimension of the Arab Spring’s impact on academia. Two other critical factors—funding and migration—also play essential roles in shaping scientific production about the region. Research funding is a key determinant of academic output, influencing the scale, focus, and sustainability of research activities  (?, ?, ?, ?, ?, ?, ?, ?). In developing countries, which often have limited domestic financial resources, the source of funding is another factor that influences research. Foreign funding plays a more pronounced role in research activities in these countries compared to those in the developed economies  (?, ?). While foreign funding enhances collaboration and citation impact, it can also lead to dependency and influence research agendas  (?, ?, ?, ?). The funding patterns observed in MENA countries exhibit significant variation, with non-MENA foreign funding assuming a substantial role in internationally co-authored publications  (?).
Scholarly migration is another fundamental consequence of the Arab Spring, significantly affecting local and global scientific landscapes. Political instability and economic hardship forced many academics to relocate, leading to a significant brain drain that weakened national research systems  (?). Although forced migration disrupts academic communities, it also creates opportunities. Displaced scholars often establish new collaborations, contribute to knowledge exchange, and drive institutional changes in their host countries  (?, ?).
@@ Introduction
In recent studies, bibliographic data has been used to track the movement of researchers across international and regional borders, revealing important migration patterns  (?, ?, ?, ?, ?, ?) and the key factors driving them  (?, ?, ?, ?, ?). Using long-term trends helps to study how major socio-political events shape brain drain and brain gain dynamics  (?), for instance, the collapse of the Soviet Union  (?, ?), Brexit  (?), the Ukraine war  (?), and China’s Cultural Revolution  (?).
Despite its importance, the Arab Spring’s impact on scholarly attention and its interplay with funding and scholarly migration remains largely unexplored. Here, we provide the first large-scale analysis of the change in scholarly attention targeted at ten countries directly impacted by the Arab Spring that underwent government overthrows (Egypt and Tunisia), civil wars (Libya, Syria, and Yemen), and governmental changes (Bahrain, Jordan, Kuwait, Morocco, and Oman). We extract country names from the titles and abstracts of more than 25 million articles and identify 73 thousand papers that mention at least one of the countries impacted by the Arab Spring. We also obtain funding information from the acknowledgments and use a scholarly migration database  (?) to understand the interplay between changes in scholarly attention, funding, and migration related to the Arab Spring. Specifically, we address the following questions:
  • How has scholarly attention toward these nations evolved compared to global trends? Are these changes uniform across countries?
  • Is the change in research activity driven more by domestic or foreign scholars? How do these dynamics vary across disciplines and countries?
  • Which countries contribute the most foreign scholarly attention, and how have their global rankings changed post-Arab Spring?
  • How do patterns of research funding—both domestic and foreign—compare to other regions? What is the relationship between funding and scholarly attention for target countries?
  • What impact has the Arab Spring had on scholarly migration? Which countries have attracted migrating scholars, and how do these patterns differ among affected nations? How does migration relate to scholarly attention?
Our results show that among the 10 target countries, Egypt is the primary focus of research, seeing an increase in scholarly attention after the Arab Spring greater than 81% of countries worldwide, followed by Tunisia. Countries in civil wars saw a surge in foreign scholarly attention, particularly in the social sciences, whereas those undergoing governmental changes experienced more varied trends. Research on the Arab Spring itself accounted for a portion of the increase but did not explain all attention shifts. While domestic scholarly attention increased across all target countries, its growth varied; Egypt and Tunisia saw stronger domestic attention, whereas the Civil War group predominantly attracted international research contributions. Foreign researchers prioritized social sciences, particularly in civil war-affected nations, whereas domestic researchers focused more on health and life sciences.
Using a difference in differences framework, we show that target countries experienced a statistically significant overall increase in scholarly attention compared to the rest of the world. In particular, the effect is positive for countries with government overthrows and civil wars. We find that the attention towards countries with governmental changes grew significantly less than towards the rest of the MENA countries. Furthermore, when breaking down the source of attention, domestic attention increased significantly in the government overthrows subgroup. Yet, the increase was less than for the rest of the world and the MENA countries in the civil war subgroup. In contrast, both these subgroups exhibited significant gains in foreign attention, while countries with governmental changes saw an increase less than the rest of the world.
Our analysis reveals that foreign-funded research significantly increased in Egypt, Tunisia, Jordan, and Morocco, primarily originating from Saudi Arabia, the United States, and the United Kingdom. The social sciences, though historically underfunded, saw a rise in financial support from 12.4% to 17.6%, likely fueling the uptick in research within this field.
@@
We show how scholarly migration further shaped the academic landscape. While Egypt remained a research hub, attracting scholars, Tunisia and Syria suffered a severe brain drain, with many researchers relocating to wealthier and more stable nations. Western and Islamic nations—especially Saudi Arabia, Qatar, the UAE, and Malaysia—emerged as top destinations for migrating researchers, presumably drawn by economic opportunities and cultural affinities. The strong correlation between scholarly migration and scholarly attention suggests that as researchers relocated, they continued producing work about their home countries, contributing to the surge in foreign attention.
@@
The complex relationship between foreign and domestic scholarly attention, funding, and scholarly migration underscores the Arab Spring’s long-term impact on academia and the power dynamics between sources and targets of scholarly attention. For example, our results reveal that Saudi Arabia leveraged its financial and academic resources to become a key player in shaping regional research, significantly strengthening bonds with Egypt, attracting their scholars, and providing funding and scholarly attention. Our study offers a foundational framework for understanding how major socio-political events shape scientific inquiry and provides insights that can apply beyond the Arab Spring to other regions undergoing major transformations.
@@ Results
The target countries in this study have all experienced some form of impact from the Arab Spring, including government overthrows, civil wars, or governmental changes, and are located in North Africa or the Middle East (Figure 1-A). These countries range from highly populous nations, such as Egypt, to less populous ones, such as Bahrain and Kuwait (see Table S1 of Supplementary Materials). Additionally, their geographical sizes vary significantly, from small nations to larger ones like Libya and Morocco.
We define scholarly attention towards a country as the extent to which it appears in the titles or abstracts of scientific publications  (?, ?). To capture the source and target of scholarly attention, we build a directed, weighted network to measure scholarly attention where each node corresponds to a country. A directed edge from country $j$ to country $i$ indicates that authors affiliated with institutions in country $j$ have mentioned country $i$ in their publication. We use fractional edge weights to account for articles with multiple authors and mentioned countries (see the Materials and Methods section for a detailed explanation). For each country $i$, the received scholarly attention in the period $\Delta t$ is the sum of incoming edge weights, i.e., $A^{i}_{tot}(\Delta t)=\sum_{j}A^{i}_{j}(\Delta t)$. The scholarly attention contributed by country $i$ towards the target countries during $\Delta t$ is the sum of outgoing edge weights to target Arab Spring countries, i.e., $A_{i}^{*}(\Delta t)=\sum_{j\in*}A^{j}_{i}(\Delta t)$, where * represents the set of target Arab Spring countries. We measure temporal changes in the network before and after the Arab Spring as the difference in yearly average scholarly attention values between two periods: 2002 to 2010 and 2011 to 2019.
@@
Figure 1: Scholarly Attention in Target Countries Pre- and Post-Arab Spring. (A) Geographical locations of the target countries. (B) Difference in total scholarly attention before and after the Arab Spring across target countries highlights non-uniform changes. In all panels, the gray area represents the interquartile range for global trends, with the dashed line indicating the median and the dotted line showing the mean. Country-specific trends reveal significant increases in the volume and distribution percentiles of scientific publications for countries experiencing government overthrow (C), civil war (D), and governmental changes (E). The vertical spans indicate periods of unrest in these countries associated with the Arab Spring and its aftermath.
@@
Figure 1-B highlights that shifts in scholarly attention, $\Delta A^{i}_{tot}$, are not evenly distributed, with certain countries drawing significantly more attention than others. While global scholarly output has generally increased over time—reflected in the upward trends of the world’s median, mean, and first and third quartiles—the extent of this growth varies among the target countries. Egypt surpasses the third quartile and the mean, whereas Tunisia, Morocco, and Jordan fall between the median and the third quartile. The remaining countries are positioned on and below the median.
The scholarly attention received by the target countries showed an increasing trend between 2002 and 2019 (Fig1-C-E). In all these countries, the annual level of scholarly attention in 2019 exceeded that of 2002. When considering the shares of the total scholarly attention allocated to each target country, we see a mostly stable trend, except for Egypt, Tunisia, and Syria, which see an increase (see Fig. S1 in Supplementary Materials).
In Egypt, scholarly attention began to rise significantly after 2009, following the global economic recession of 2008, exceeding the global third-quartile trend, and from 2011 onward, it rose above the worldwide average and remained there until 2015 (Fig 1-C). Although there was a decline in 2016, the trend began to rise again, stabilizing around the global mean. Tunisia experienced a notable rise starting in 2005, which elevated it above the global median level. This growth plateaued after 2010, with scholarly attention remaining relatively stable since then rather than continuing to increase.
Among the countries with civil wars, Yemen and Libya display similar patterns in the level of scholarly attention they receive, which is different from Syria’s (Fig 1-D). Before 2008, research on these countries was minimal, which placed them around the first quartile regarding global distribution. A significant increase occurred in 2009, followed by a plateau that lasted until the onset of the Arab Spring. In 2012, a year after the Arab Spring, another increase in attention was observed. Libya peaked in scholarly attention in 2013, while Yemen peaked in 2014. After this peak, both countries experienced a relatively consistent decline, eventually falling below the first quartile. Syria experienced a mostly upward trend, with a notable peak in 2016 (probably related to the rise of ISIS and the humanitarian crisis it caused), consistently elevating Syria closer to the median, reflecting increased global research interest.
Countries that experienced governmental changes exhibit diverse trends in scholarly attention (Fig 1-E). Jordan and Morocco show a steady upward trend, consistently positioning them between the third quartile and median. Unlike other countries that saw a significant increase during the unrest period, Oman experienced a decline, from which it recovered, returning to near the median. Kuwait stayed relatively stable, with minimal fluctuations, gradually shifting to lower percentiles. Bahrain presents a different pattern with limited research focused on it. Its trend features two distinct peaks: one in 2008, followed by a decline in 2009, and another in 2013, which declined shortly afterward. After 2015, Bahrain’s scholarly attention began a gradual upward trajectory, eventually bringing it closer to the first quartile.
@@
To understand what part of the increase in scholarly attention is due to research directly focusing on the Arab Spring, we investigate articles containing mentions of keywords related to the “Arab Spring” (e.g., Arab Uprisings, Arab Revolution, etc., see the Materials and Methods section for a comprehensive list and explanation of the methodology) in their title or abstract. In Figure 2, we show that only a fraction of the scholarly attention to a target country following the Arab Spring can be credited to research related to the Arab Spring.
Figure 2: Scholarly attention due to Arab Spring-related research. (A) Country-specific scholarly attention related to research on the Arab Spring. Research focusing on the government overthrow (GO) group attracted the highest scholarly attention, followed by the civil war group (CW). (B) Proportion of Arab Spring-related research as part of total scholarly attention by target country. Syria and Libya, classified under the civil war group, exhibit the highest shares, whereas the governmental changes group shows minimal shares. The vertical dotted line represents the year 2011, marking the onset of the Arab Spring.
The government-overthrown group attracted the most attention in Arab Spring-related research, followed by the civil war group. Egypt also attracted the most attention (Fig 2-A). The proportion of scholarly attention related to the Arab Spring, compared to the total attention received, reveals that the civil war group holds a significantly larger share than others. Libya recorded the highest share in 2013, accounting for almost 14%. Syria shows two notable peaks, one in 2012 and another in 2014. Interestingly, Bahrain ranks among the top three in share but experienced a decline after 2014. The government overthrow group demonstrates relatively steady attention, maintaining a consistent share of around 3-4%. For the remaining countries, the share is generally below 2% of the total.
@@
Scholarly attention to a country can come from two main sources: researchers affiliated with that country (domestic attention), namely $A_{i}^{i}$ (self-loops in the scholarly attention network), and researchers from other countries (foreign attention), namely $\sum_{j\neq i}A_{j}^{i}$.
@@
Figure 3-A shows that following the Arab Spring, domestic attention rose across all target countries, although the extent of this increase varied. While civil war and governmental change groups showed only modest gains, the government-overthrown group sustained significant domestic attention, possibly due to a robust local scholarly community and strong, established universities. Similarly, while foreign scholarly attention increased for all three groups (Fig 3-B), the increases in the governmental changes group are significant, more than the rise for the world’s median. Syria, notably, transitioned from being below the median before the Arab Spring to above it afterward.
The foreign-to-domestic attention ratio (Fig. 3-C) quantifies the relative contribution of each source for one country. In Egypt, the ratio dropped to around 50%, indicating that domestic attention grew more than foreign attention. Oman and Morocco transitioned from ratios above 1 to around and below 1, respectively, possibly because of growth in domestic research funding or productivity, which boosted domestic attention beyond foreign attention. Libya, Syria, and Yemen had ratios above 1 before the Arab Spring, increasing further afterward, highlighting growing foreign attention.
@@
Figure 4: Share of domestic and foreign scholarly attention before and after the Arab Sp ring by discipline. The target countries are split into three groups: government overthrow (GO), civil war (CW), and governmental changes (GC). Filled bars represent values before the Arab Spring, while hatched bars indicate values after the Arab Spring. The world’s median before and after the Arab Spring are marked with solid and dashed black lines, respectively. In domestic attention (A–D), social sciences consistently held the smallest share for most countries. After the Arab Spring, the proportion of social sciences in the governmental changes group (excluding Morocco) rose above the median. In foreign attention (E–H), social sciences showed a notably higher share—particularly in Yemen and Libya, where it became the dominant field post-Arab Spring.
@@
To understand how different disciplines resonate with domestic and international audiences, we analyze foreign and domestic scholarly attention across four major disciplines: social sciences, health sciences, life sciences, and physical sciences (see Fig. 4). In domestic attention (Figure 4-A-D), the relative shares of each discipline show that in countries belonging to the government overthrow and civil war groups, as well as Morocco, the social sciences consistently held the smallest share of attention both before and after the Arab Spring. Although the other groups did not show notable shifts, after the Arab Spring, the proportion of social sciences in the government changes group (excluding Morocco) rose above the median. This shift could be attributed to government reforms, which may have reallocated attention toward social sciences. We also note that scholarly attention to health sciences in Yemen declined significantly following the Arab Spring.
In foreign attention (Fig. 4-E-H), the social sciences consistently exhibit a much higher proportion than what is observed in domestic attention. This divergence is mainly observed in the civil war group, such as in Yemen and Libya, where social sciences, once below the median before the Arab Spring, rose above the median afterward and became the dominant discipline. Notably, their share increased by 21.5% in Syria and by 29.5% in Libya, suggesting increased international interest in understanding social and political dynamics in these conflict-affected nations. By contrast, in Morocco and Oman, the physical sciences receive the largest share of foreign attention, possibly driven by the regions’ historical and geographical research appeal. In Libya, the share of foreign attention to physical sciences decreased significantly following the Arab Spring.
@@
Figure 5: Geographical distribution of foreign scholarly attention sources towards target countries: (A) Differences in average yearly scholarly attention towards all target countries: western countries alongside Saudi Arabia show the highest differences, while countries in the Global South show the least. (B) Rank inertia capturing relative attention changes: The impact of the Arab Spring is more evident in the region and among Islamic countries. Saudi Arabia is present in the top 5 countries in terms of the difference in average and rank inertia metrics, highlighting its significant influence in research on the target countries. Territories are visualized with Geopandas  (?).
Main sources of the increase in scholarly attention

### v3

@@
Rezvaneh Rezapour6, Mary Ellen Sloane7, and Alexandre Bovet 1,2∗
@@
University of Zurich, 8057 Zurich, Switzerland
@@
{Yasaman.asgari, Alexandre.bovet}@uzh.ch
The 2010–2011 Arab Spring reverberated far beyond politics, reshaping how the Middle East and North Africa region (MENA) is studied. Analyzing 3.7 million Scopus-indexed articles published between 2002 and 2019, we find that mentions of ten of these countries in titles or abstracts rose significantly after 2011 relative to the global baseline, with Egypt receiving the greatest attention in the region. We link this surge to two intertwined mechanisms: an increase in research funding directed at the MENA region and the emigration of researchers who continued publishing on their countries of origin. Our analysis reveals that Saudi Arabia has emerged as a regional hub for studying the affected countries, attracting funding and scholars, and thereby playing a significant role in shaping the scientific narrative on the region. These findings demonstrate how political upheaval can reshape global knowledge flows by altering who studies whom, with what resources, and in which disciplines.
Main
Science and research do not happen on their own; they are shaped by and help shape the economic and sociopolitical environment of their societies  (?, ?, ?). Major events, such as revolutions, armed conflicts, and natural disasters, profoundly disrupt scientific ecosystems, resulting in reduced research productivity, weakened international collaboration, and increased brain drain  (?, ?, ?, ?, ?, ?). The Arab Spring, sparked by Mohamed Bouazizi’s self-immolation in Tunisia in 2010, exemplifies this phenomenon across the Middle East and North Africa (MENA), affecting nearly every country in the region  (?).
To date, investigations into the Arab Spring’s effects on scholarly ecosystems have relied mainly on bibliometric analysis with a focus on collaboration patterns  (?, ?, ?), academic freedom  (?, ?), and research productivity  (?, ?, ?). More recent works in political science, international relations, and sociology have shown a marked surge in academic discourse on the transformations of the Middle East and North Africa (MENA) following the uprisings  (?, ?).
However, there remains a notable absence of large-scale, systematic analyses that (1) identify who is studying which MENA countries, (2) trace how this attention varies across disciplines, and (3) reveal the evolution of this attention over time. Here, we address this gap by defining scholarly attention as the frequency with which countries are mentioned in article titles and abstracts, and take this as a first lens to understand the effect of the Arab Spring on the geographic distribution of research focus. Country naming practices have already proven effective in capturing global research patterns in domains ranging from public health  (?) and the social sciences  (?) to research on the United Nations’ Sustainable Development Goals (SDGs)  (?).
A second lens is research funding: naturally, research output is primarily a function of the funding available; thus, it is essential to track the origins of funding. Indeed, funding plays a decisive role in determining the scale, thematic focus, and sustainability of research activities  (?, ?, ?, ?, ?, ?, ?, ?). In developing countries, which often have limited domestic financial resources, foreign funding plays a more pronounced role in research activities compared to those in the developed economies  (?, ?, ?). While foreign funding enhances collaboration and citation impact, it can also create dependency and influence on research agendas  (?, ?, ?, ?).
@@
A third lens is scholarly migration: when researchers relocate, they often continue research on their home country, sustaining or redirecting attention from abroad. Bibliometric data have been recently used to track the movement of researchers across borders, revealing important migration patterns  (?, ?, ?, ?, ?, ?) and the key factors driving them  (?, ?, ?, ?, ?). Using long-term trends helps scholars to study how major socio-political events shape brain drain and brain gain dynamics, for instance during World War II  (?), in China’s Cultural Revolution  (?), the collapse of the Soviet Union  (?, ?), Brexit  (?), and the Ukraine war  (?).
Together, these three lenses—attention, funding, and migration—offer a holistic view of how the Arab Spring reshaped global research dynamics. Moreover, examining the link between attention and both funding flows and scholarly migration gives a unique perspective on the forces shaping these shifts. Here, we present a longitudinal analysis of scholarly attention towards ten countries affected by the Arab Spring, drawing on Ibrahim’s 2018 classification of sociopolitical trajectories into government overthrows (Egypt, Tunisia), civil wars (Libya, Syria, Yemen), and governmental changes (Bahrain, Jordan, Kuwait, Morocco, Oman), as well as major and minor protests (?).
We extracted mentions of countries from the titles and abstracts of over 25 million articles published between 2002 and 2019 and identified 73,000 papers that mention at least one of the target countries. As a comparison baseline, we also analyzed 3.7 million articles that mention any country. Moreover, we obtained funding information embedded in acknowledgments and used a scholarly migration database  (?) to examine the relationship between changes in scholarly attention, funding, and migration. We employed a difference-in-differences approach to estimate the causal impact of the Arab Spring on scholarly attention.
Our findings indicate that the target countries experienced a statistically significant increase in scholarly attention compared to the rest of the world. Within this group, Egypt emerged as the primary focus of research, with its post-uprising attention growth exceeding 80% of all the countries globally. Attention patterns differed by researcher affiliation: foreign scholars disproportionately contributed to social science publications, whereas domestic scholars focused more on the health and life sciences. Foreign-funded research increased significantly, primarily from Saudi Arabia, the United States, and the United Kingdom. The social sciences, although historically underfunded, have seen increased funding, which has fueled research growth in this field.
@@
Regarding scholarly migration, Egypt remained a research hub, attracting scholars, while Tunisia and Syria experienced severe brain drain as researchers relocated to wealthier nations. Western and Islamic countries, especially Saudi Arabia, Qatar, the UAE, and Malaysia, emerged as top destinations, presumably due to economic opportunities and cultural affinities. The strong correlation between scholarly migration and scholarly attention suggests that relocated researchers continued studying their home countries, contributing to increased foreign attention.
@@
These findings underscore the Arab Spring’s lasting impact on academia and reveal the power dynamics between sources and targets of scholarly attention. The relationship between scholarly attention, funding, and scholarly migration demonstrates how major socio-political events reshape scientific inquiry. Our study provides a foundational framework for understanding these dynamics, offering insights applicable to other regions undergoing major transformations.
@@ Results
The target countries in this study are located in North Africa or the Middle East (Figure 1a). These countries range from highly populous nations, such as Egypt, to less populous ones, including Bahrain and Kuwait (see Table S1 in the Supplementary Materials).
We quantify the scholarly attention these countries received via their mentions in Scopus publication titles and abstracts, which we represent as a directed, weighted network: each node corresponds to a country, and a directed edge from country $j$ to country $i$ fractionally counted number of papers authored by researchers at institutions in $j$ that mention $i$ (see the Materials and Methods section for a detailed explanation). A country’s total received attention is given by $A^{i}_{\mathrm{tot}}=\sum_{j}A^{i}_{j}$, and temporal change in attention is computed as the difference in its average annual attention between the pre-Arab Spring period (2002–2010) and the post-Arab Spring period (2011–2019).
@@
Figure 1: Scholarly Attention in Target Countries Pre- and Post-Arab Spring. (a), Geographical locations of the target countries. (b), The difference in total scholarly attention before and after the Arab Spring across target countries highlights non-uniform changes. In all panels, the gray area represents the interquartile range for global trends, with the dashed line indicating the median and the dotted line showing the mean. Country-specific trends reveal significant increases in the volume and distribution percentiles of scientific publications for countries experiencing government overthrow (c), civil war (d), and governmental changes (e). The vertical spans indicate periods of unrest in these countries associated with the Arab Spring and its aftermath.
@@
Figure 1b highlights that shifts in scholarly attention towards each country $i$, namely $\Delta A^{i}_{tot}$, are not evenly distributed across the countries. Egypt surpasses the third quartile and the mean, whereas Tunisia, Morocco, and Jordan fall between the median and the third quartile. Meanwhile, the remaining countries are positioned on and below the median.
While the attention temporal trajectories differ across these countries, all of them received more annual scholarly attention in 2019 than in 2002 (Fig1c–e). Research on Egypt began accelerating after 2009, then rose above the global average between 2011 and 2015, dipped slightly in 2016, and then stabilized around the world’s mean (Fig. 1c). Attention towards Tunisia climbed steadily from 2005 until about 2010, after which it hovered at a stable plateau. In the civil war group (Fig. 1d), Yemen and Libya received minimal attention until 2009, followed by a surge that plateaued until 2012; Libya reached its peak in attention in 2013, while attention towards Yemen peaked in 2014, before both declined. Syria experienced a steady rise in attention, with a peak in 2016 probably corresponding to the rise of ISIS and associated humanitarian crises. Among the governmental changes group, attention towards Jordan and Morocco rose above the global median consistently; attention directed at Oman experienced a brief decline in 2012; attention towards Kuwait remained relatively flat; and Bahrain, despite a low overall volume, showed peaks in attention in 2008 and 2013, followed by growth after 2015.
When normalizing each country’s attention relative to the total global attention, the fraction of attention targeted at most countries displays stable trajectories, except for Egypt, Tunisia, and Syria, which show notable increases (see Fig. S1 in the Supplementary Materials).
@@
One possible explanation for the rise in scholarly attention is the increased research related to the Arab Spring. To assess how much of this attention is specifically driven by the Arab Spring’s aftermath, we investigate articles containing mentions of keywords related to the “Arab Spring” (e.g., Arab Uprisings, Arab Revolution, etc., see the Materials and Methods section for a comprehensive list and explanation of the methodology) in their title or abstract.
Figure 2: Scholarly attention due to Arab Spring-related research. (a), Country-specific scholarly attention related to research on the Arab Spring. Research focusing on the government overthrow (GO) group attracted the highest scholarly attention, followed by the civil war group (CW). (b), Proportion of Arab Spring-related research as part of total scholarly attention by target country. Syria and Libya, classified under the civil war group, exhibit the highest shares, whereas the governmental changes group shows minimal shares. The vertical dotted line represents the year 2011, marking the onset of the Arab Spring.
In Figure 2, we show that only a small fraction of the scholarly attention to a target country following the Arab Spring can be credited to Arab Spring-related research. The government-overthrown group, specifically Egypt, attracted the most attention in Arab Spring-related research, followed by countries in the civil war group (Fig. 2a). The proportion of scholarly attention related to the Arab Spring, compared to the total attention received, reveals that the civil war group holds a significantly larger share than the others(Fig. 2b). The government overthrow group demonstrates relatively steady attention share around 3-4%, and for the governmental changes group, except Bahrain, the share is generally below 2%. Libya recorded the highest share in 2013, accounting for almost 14%. Syria shows two notable peaks, one in 2012 and another in 2014. Bahrain’s presence among the top three in terms of share is due to its low total attention.
@@
We now turn to identifying the sources of scholarly attention toward the target countries. Scholarly attention to a country can come from two sources: researchers affiliated with that country (domestic attention), namely $A_{i}^{i}$ (self-loops in the scholarly attention network), and researchers from other countries (foreign attention), namely $\sum_{j\neq i}A_{j}^{i}$.
@@
Figure 3a shows that domestic scholarly attention increased across all target countries following the uprisings, though with substantial variation. The largest increase occurred in the government-overthrown group, which consistently received a higher level of domestic attention. This increase may reflect a revitalized local research environment supported by established universities and national rebuilding efforts. Meanwhile, countries in the civil war group showed only modest gains in domestic attention, a pattern likely influenced by prolonged conflict, institutional disruption, reduced research capacity, and large-scale scholarly migration. The governmental change group also experienced notable increases, particularly in Jordan and Morocco, where internal reforms may have contributed to a more stable environment for domestic scholarly activity.
@@
Similarly, while foreign scholarly attention increased for all three groups (Fig. 3b), the increase in the government overthrow group is significant, surpassing the rise for the world’s median. In the civil war group, Syria, notably, transitioned from being below the median before the Arab Spring to above it afterward.
The foreign-to-domestic attention ratio (Fig. 3c) quantifies the relative contribution of foreign versus domestic scholarly attention for each country. In Egypt, the ratio declined to approximately 0.5 after the Arab Spring, indicating that domestic attention outpaced foreign attention. Oman and Morocco both transitioned from foreign-dominated attention (ratios above 1) to more balanced ratios, close to one. This shift may reflect an increase in domestic research capacity or institutional support. By contrast, the civil war group had ratios above one before the uprisings, which rose even further afterward, signaling a growing dominance of foreign over domestic attention.
These patterns show that domestic and foreign scholarly attention did not evolve uniformly across cases. Countries with more stable political transitions saw relatively greater domestic engagement, while those experiencing sustained conflict attracted increasing levels of international attention. Higher foreign-to-domestic attention ratios may be associated with international concern over conflict spillovers or humanitarian crises.
@@
Figure 4: Share of domestic and foreign scholarly attention before and after the Arab Spring by discipline. The target countries are split into three groups: government overthrow (GO), civil war (CW), and governmental changes (GC). Filled bars represent values before the Arab Spring, while hatched bars indicate values after the Arab Spring. The world’s median before and after the Arab Spring is marked with solid and dashed black lines, respectively. In domestic attention (a–d), social sciences consistently held the smallest share for most countries. After the Arab Spring, the proportion of social sciences in the governmental changes group (excluding Morocco) rose above the median. In foreign attention (e–h), social sciences showed a notably higher share, particularly in Yemen and Libya, where it became the dominant field post-Arab Spring.
@@
Figure 4 shows how foreign and domestic attention differ across four major disciplines: social sciences, health sciences, life sciences, and physical sciences. In domestic attention (Fig. 4a-d), the relative shares of each discipline show that in the government overthrow and civil war groups, as well as Morocco, the social sciences consistently held the smallest share of attention before and after the Arab Spring. Although the other groups did not show notable shifts, after the Arab Spring, the proportion of social sciences in the government changes group (excluding Morocco) rose above the median. This shift could be attributed to government reforms, which may have reallocated attention toward social sciences. Meanwhile, the relative share of domestic scholarly attention to health sciences in Yemen has declined significantly following the Arab Spring, due to an increase in domestic attention to other disciplines (for analysis of the counts, refer to Supplementary Materials, Fig. S2). In foreign attention (Fig. 4e-h), the social sciences consistently exhibit a much higher proportion than what is observed in domestic attention. In Yemen and Libya, the social sciences, which had previously been below the median before the Arab Spring, rose above the median afterward and became the dominant discipline. Moreover, in Morocco and Oman, the physical sciences receive the largest share of foreign attention, possibly driven by the regions’ historical and geographical research appeal.
We also computed the foreign-to-domestic attention ratio across individual disciplines (see Fig. S3 in the Supplementary Material). The decrease in this ratio for Egypt occurs across all disciplines, indicating a consistent pattern of increased domestic attention following the Arab Spring. In the civil war group, the social sciences have extremely high foreign-to-domestic ratios, highlighting a heavy reliance on foreign sources. Similarly, in the health sciences, foreign-to-domestic ratios rose, shifting from below one to above one. In the governmental changes group, except for Kuwait, the foreign-to-domestic attention ratio in social sciences decreased, showing a greater increase in domestic attention. These diverse patterns suggest a strong dependence of scholarly attention on political trajectories.
Main sources of the foreign increase in scholarly attention
@@
Figure 5: Geographical distribution of foreign scholarly attention sources towards target countries: (a) Differences in average yearly scholarly attention towards all target countries: Western countries alongside Saudi Arabia show the highest differences, while countries in the Global South show the least. (b) Normalized rank change capturing relative attention changes: The impact of the Arab Spring is more evident in the region and among Islamic countries. Saudi Arabia is among the top 5 countries in terms of the difference in average and rank inertia metrics, highlighting its significant influence in research on the target countries. Territories are visualized with Geopandas  (?).
We now examine the geographic sources of foreign scholarly attention (see Figs. 5 and 6) to identify which regions and countries contributed to the observed increase in foreign attention toward Arab Spring-affected nations.